OpenStudy (anonymous):

Another Calc II question. Find y'(x) if y= integral of sin(2e^t) dt *limits of integration are from 0 to ln(4x)* I don't understand the y'(x) part really.

OpenStudy (amistre64):

\[\frac d{dx}~(\int_{a(x)}^{b(x)}f(x)~dx)=f(b)b'-f(a)a'\]

OpenStudy (amistre64):

we can follow this using the definitions of integration and derivatives

OpenStudy (amistre64):

\[\int_{a(x)}^{b(x)}f(t)dt=F(b(x))-F(a(x))\] take the derivative of that, noting the chain rule \[Dx[F(b(x))-F(a(x))]=f(b(x))b'-f(a(x))a'\]

OpenStudy (anonymous):

I understand how to integrate but why does it say y'(x) because to integrate you are finding the original function.

OpenStudy (amistre64):

yes, notice that the integration creates a function of "x"

OpenStudy (anonymous):

but when i integrate as well I cannot use the U substitution because setting u=(2e^t) would be du/dx = 2e^t and there isn't anywhere in the integral containing that

OpenStudy (amistre64):

notice the form from the definitions; dont try to do all the work .. you cant

OpenStudy (amistre64):

what is f(t) ?

OpenStudy (anonymous):

-cos(2e^t)?

OpenStudy (amistre64):

according to the post: f(t) = sin(2e^t)

OpenStudy (anonymous):

ohhh im so confused!

OpenStudy (amistre64):

what then is f(b(x)) ? and f(a(x)) ?

OpenStudy (anonymous):

sin(2e^ln(4x))-sin(2e^0)?

OpenStudy (amistre64):

good, now lets fill in the missing parts; the b' and a' sin(2e^ln(4x)) * (ln(4x))' - sin(2e^0)*0'

OpenStudy (amistre64):

the key is to know that by definitions of integration and derivative, we have all the parts we need. It doesnt matter what the solution of the integration is, becasue we end up taking the derivative back down again

OpenStudy (anonymous):

sin(2e^ln(4x)) * (1/4x)?

OpenStudy (amistre64):

close, your ln(4x) derivative needs work i think

OpenStudy (amistre64):

\[D_x(ln(u))=\frac{u'}{u}\] \[D_x(ln(4x))=\frac{4}{4x}\to\ \frac{1}{x}\]

OpenStudy (anonymous):

right i just did that can and can i simplify sin(2e^ln(4x))

OpenStudy (amistre64):

im sure you can \[e^{ln(4x)}=e^{ln4+lnx}=e^{ln4}*e^{lnx}\to\ 4x\]

OpenStudy (anonymous):

right i got sin(8x) (1/x)

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

and then i just derive it from there?

OpenStudy (amistre64):

derive what? all youre doing is simplifying what youve already derived

OpenStudy (anonymous):

so that's y'(x)?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

okay...im still a bit confused with that though

OpenStudy (anonymous):

i understand the process but the concept is a bit confusing

OpenStudy (amistre64):

\[y(x)=\int_{a(x)}^{b(x)}f(t)~dt\] \[y(x)=F(b(x))-F(a(x))\] you agree with this by definition of integration?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then y'(x) is the derivative\[y'(x)=[F(b(x))−F(a(x))]'\] then y'(x) is the derivative\[y'(x)=[F(b(x))]'−[F(a(x))]'\] then using the chain rule; y'(x) is the derivative\[y'(x)=F'(b(x))b'−F'(a(x))a'\] agreed?

OpenStudy (anonymous):

ohh yess i see now

OpenStudy (amistre64):

and we already now what F' is.... its just the f that we started with

OpenStudy (anonymous):

koayy thanks you so much I understand now. Much clearer

OpenStudy (amistre64):

youre welcome :)